Hello everyone
Please help my granddaughter with this problem
Video games cost $14.50 each, shipping is flat rate of $6, you have $50 to spend, how many games can you buy?
Thank you
Gina

Answer by jorel1380(3719)

6+14.50x=50
14.50x=44
x=44/14.50=3.0345
You can buy 3 games with your money..